OpenStudy (anonymous):

Given the function f described by f(x)=x+12, find... f(0)= f(-5)= f(-12)= f(8)= f(x+16)=

OpenStudy (anonymous):

f(-12)

OpenStudy (anonymous):

So you just replace the X with what is in the parenthesis

OpenStudy (anonymous):

so f(0)=x+12 i don't understand this problem

OpenStudy (anonymous):

replace x with 0 so f(0) =12

OpenStudy (anonymous):

You just put in whatever is in parens for the x. f(0) = 0 + 12 = 12

OpenStudy (anonymous):

f(8) = 8 + 12 = 20

OpenStudy (anonymous):

etc.

OpenStudy (anonymous):

ok pol so i figured those out, how do i do f(x+16)=?

OpenStudy (anonymous):

f(x+16)=? replace x in f(x)=x+12 by (x+16) f(x+16)= x+16+12 = x +28